OpenStudy (anonymous):

Find the slope of the line that contains the following pair of points. (1, 2), (5, 2) A. -1 B. 0 C. 4 D. Undefined

OpenStudy (aravindg):

\[\large slope=\dfrac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Slope is 0 Point 1 (1,2) Point 2 (5,2) x1=1 , x2 = 5 and y1 = 2 , y2 = 2 (2-2)/(5-1) 0 / 4 = 0

OpenStudy (anonymous):

check my work please?

OpenStudy (aravindg):

right !

OpenStudy (anonymous):

Sweet! Thank you :)

OpenStudy (diyaja):

slope is (2-2)/5-1). the answer is undefined

OpenStudy (aravindg):

infact think of it graphically they have same y value means it will be a horizontal line so slope=0

OpenStudy (aravindg):

@DiyaJA slope =0

OpenStudy (anonymous):

the equation is (2-2)/(5-1)